[Alert_collaboration] [New Logentry] RC Daily Summary
sjjooste at jlab.org
sjjooste at jlab.org
Mon May 12 10:00:01 EDT 2025
Logentry Text:
--
<div id="content"><h1 id="rc-summary-for-2025-05-12---mon-may-12">RC Summary for 2025-05-12 - Mon May 12</h1>
<p><strong>We will have the RC meeting at 11:30am EDT in the Counting House. <a href="https://jlab-org.zoomgov.com/j/1605705819#success">Zoom Link</a></strong></p>
<h2 id="recap-and-status">Recap and Status</h2>
<h4 id="sun-may-11-dayhttpslogbooksjlaborgentry4373085--swinghttpslogbooksjlaborgentry4373286---mon-may-12-owlhttpslogbooksjlaborgentry4373386">Sun May 11 <a href="https://logbooks.jlab.org/entry/4373085">DAY</a> & <a href="https://logbooks.jlab.org/entry/4373286">SWING</a> - Mon May 12 <a href="https://logbooks.jlab.org/entry/4373386">OWL</a></h4>
<ul>
<li>No beam DAY or SWING</li>
<li>Riad Suleiman from ACC (injector) came to the counting house, and was asking if we have a way to check if there is a time structure (10-50ms ) in our halo rates --> <a href="https://logbooks.jlab.org/comment/53020#comment-53020">Nathan made the 60Hz display working</a>, which gives us an FFT</li>
<li><a href="https://logbooks.jlab.org/entry/4373122">Lost EPICS communication</a> due to the Hall B <a href="https://logbooks.jlab.org/entry/4373125">EPICS gateway going down</a> during around 10:30am . Alarm server still had proper connections, and Hall B internal IOCs were still accessible. We could also access external EPICS IOCs through the Hall A/Hall C gateways. MCC looped in tech support, who (eventually) looped in an expert. Expert immediately diagnosed it as an improper gateway reboot and fixed it by killing a rogue process on the gateway. It was resolved around 1:30pm. This was not a safety issue (internal alarms/controls still worked), and we did not lose beam (as there was no beam to lose).</li>
<li>During SWING we saw <a href="https://logbooks.jlab.org/entry/4373300">high bleed-through rates</a> during various tests</li>
<li>MCC determined the biggest impact on the beam halo rates is the horizontal (!) separator. As one of the Arc-A power supplies is down anyway, preventing beam to Hall D, they made the call to magnetically steer the beam to the BSY, and then do regular RF horizontal separation into A-B-C. They started beam restoration to Hall A around midnight.</li>
<li>We started beam restoration in OWL around 3am. Tagger dump harp scans looked good. When we moved to the faraday cup, we saw very high BOM rates, and large beam offsets (6-7mm off in y). We contacted Eugene for guidance on how to proceed, and he worked directly with the operator to fix the steering issue - took about 2 hours until we were ready to do harp scans. The high BOM rates were caused by the beam going <em>outside</em> the BOM, and we had to take a larger jump to go inside without passing <em>through</em> the BOM which would trip the beam. Harp scans looked a bit wide so we had them tighten the beam. By 8am we had a nice beam profile and starting a transition to production</li>
<li>Beam halo rates looking good!</li>
</ul>
<h4 id="current-status">Current Status</h4>
<ul>
<li>During start of DAY, We did some running on 30nA with empty and He-4 target, and had MCC address some slight beam misalignment issues. We are currently running production on He-4 and preparing for ALERT repairs</li>
<li>There is NO scheduled maintenance day this week. However, they are currently working on a plan for the week as they have to:<ul>
<li>Fix the Arc-A power supply (which will take beam down)</li>
<li>Address the beam quality issue when both 5th pass separators are running</li>
<li>It's quite likely they will have substantial downtime this week.</li></ul></li>
</ul>
<h2 id="game-plan">Game Plan</h2>
<ul>
<li><a href="https://wiki.jlab.org/clas12-run/index.php/Run_Group_L#tab=Short_Term_Schedule">Short Term Schedule</a></li>
<li>Production runs with He-4 until Wednesday, aim for 100M events<ul>
<li>Monitor BOM scalers closely</li>
<li>Perform Hall A bleed-through test for ~5 minutes once Hall A is in stable running conditions with 20µA beam</li></ul></li>
<li>~Wednesday start ALERT repairs (note: NO scheduled maintenance)<ul>
<li>Opportunistically start preparing in the Hall on Tuesday if there is an early down</li></ul></li>
</ul>
<h2 id="bta">BTA</h2>
<ul>
<li>327.1 ABU hours since April 5th. 327.1/(68.5 x 24) = 19.9%</li>
<li>We are on day 37 of 137 scheduled days: 35/137 = 27.0%</li>
<li>BANU: 0.6 (beam restoration)</li>
</ul>
<h2 id="outstanding-major-tasks">Outstanding Major Tasks</h2>
<ul>
<li>Repairs planned for Wednesday, May 14th:<ul>
<li>ATOF module 7 (Whitney Armstrong)</li>
<li>AHDC wire repair (Raphael Dupre)</li>
<li>Draft of repair plans is available to all relevant people to make sure the repairs follow proper procedures and to ensure smooth execution and minimal downtime.</li></ul></li>
<li>Faraday Cup Vacuum to be monitored, to be fixed after end-of-run</li>
</ul>
<h2 id="run-summary">Run Summary</h2>
<p>N/A (no beam)</p>
<h2 id="links">Links</h2>
<ul>
<li><a href="https://jlab-org.zoomgov.com/j/1605705819#success">RC Meeting Zoom Link</a> (no RC meeting Sa-Su)</li>
<li><a href="https://wiki.jlab.org/clas12-run/index.php/Run_Group_L">ALERT/RG-L Wiki</a></li>
<li><a href="https://wiki.jlab.org/clas12-run/index.php/Run_Group_L#:~:text=General-,Logbook,-Shift%20Schedule">Hall B Logbook</a></li>
<li><a href="https://wiki.jlab.org/clas12-run/index.php/Run_Group_L#tab=Short_Term_Schedule">Short Term Schedule</a></li>
</ul></div>
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4373757
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/alert_collaboration/attachments/20250512/96be0abb/attachment.htm>
More information about the Alert_collaboration
mailing list